# bibliography.bib

### v3

@@ 3.3. General linear groups and subgroups
where $\underline{\operatorname{End}}$ is an internal Hom. Hence $\mathfrak{gl}(m+nP)$ is an $(m+2n)^{2}$-dimensional vector space, and $d$ acts via $1\otimes d+d\otimes 1$. The bracket operation is given by $[\phi_{1},\phi_{2}]=\phi_{1}\phi_{2}-\phi_{2}\phi_{1}+\phi_{2}^{\prime}\phi_{1}^{\prime}$.
@@ 4.2.1. Representations of $\operatorname{GL}(P)$
$0\to\wedge^{3}T_{1}\to T_{1}\otimes\wedge^{2}T_{1}\to S^{(2,1)}T_{1}\to 0$
@@
For degree $4k$, note that we have a non-split exact sequence $0\rightarrow\chi^{k}\rightarrow T_{4k}\rightarrow\xi\chi^{k}\rightarrow 0$. For degree $14k+$ 1, note that $T_{4}$ is a nontrivial extension of $\chi$ by $\xi$, so $\chi^{k-1}T_{4}T_{1}$ is a nontrivial extension of $\chi^{k}T_{1}$ by $\xi\chi^{k}T_{1}$. For degree $34k+$ 3, we showed above that $T_{1}T_{2}$ is a nontrivial extension of $T_{3}$ by $\xi T_{3}$, so $\chi^{k}T_{1}T_{2}$ is a nontrivial extension of $\chi^{k}T_{3}$ by $\xi\chi^{k}\otimes T_{3}$. ∎
@@
Recall that $\mathcal{O}(M_{1})=k[Y,Y^{\prime}]$ with $\Delta(Y)=Y\otimes 1+(1+Y^{\prime})\otimes Y$. Then we can compute $\Delta(Y^{n})$ and $\Delta(Y^{n}Y^{\prime})$ by using that $Y^{4\cdot 2^{k}}$ is primitive and computing $\Delta(Y^{n})$ for $30\leq n<$ 3, e.g. $\Delta(Y^{2})=1\otimes Y^{2}+Y^{2}\otimes 1+Y^{\prime}\otimes Y^{\prime}$. Note the $Y^{\prime}\otimes Y^{\prime}$ term appears because the braiding is nontrivial, so
@@ 4.2.2. Explicit structure of representations of $\operatorname{GL}(\mathds{1}+P)$ and their tensor products
Let $L(n,T)$ be the irreducible representation of $\operatorname{GL}(1+P)$ corresponding to $n$ an integer (i.e. a representation of $\operatorname{GL}_{1}$) and $T$ an irreducible representation of $\operatorname{GL}(P)$. Because $L(2,\mathds{1})$, $L(0,\chi)$ are one-dimensional, it suffices to describe $L(n,T)$ when $n=0,1$ and $40\leq\deg T<$ 4.
@@
The weight decompositions for $\operatorname{GL}(\mathds{1}+P)$ representations with highest weight $(n,T)$, $0<\deg T\leq 4$, $20\leq n<$ 2, are listed below. Extensions are denoted by $\times$.
@@ 4.2.3. Representations of $\operatorname{GL}(nP)$
The augmentation ideal $I\subset\mathcal{O}(\operatorname{GL}(nP))$ is generated by $C_{ij}-\delta_{ij}$ where $\delta$ is the Kronecker delta, and $D_{ij},C^{\prime}_{ij},D^{\prime}_{ij}$ for $1\leq i,j\leq n$. For ease of notation, let $C-1:=\{C_{ij}-\delta_{ij}\}$, $D:=\{D_{ij}\}$, $C^{\prime}:=\{C^{\prime}_{ij}\}$, $D^{\prime}:=\{D^{\prime}_{ij}\}$. Then, the residue classes of monomials in $C-1$, $D$, $C^{\prime}$, $D^{\prime}$ of degree at most $m$ form a basis for $\mathcal{O}(G)/I^{m+1}$ and for a fixed monomial $p(C-1,D,C^{\prime},D^{\prime})$, there is a unique $\delta_{p}\in\operatorname{Dist}(G)$ such that
@@
when $q\neq p$ is another monomial in $C-1,D,C^{\prime},D^{\prime}$ of degree at most $m$. The $\delta_{p}$ over all monomials $p$ of any nonnegative degree form a basis of $\operatorname{Dist}(G)$. We claim that as a $\operatorname{Dist}(G_{*})$-module, $\operatorname{Dist}(G)$ is generated by $\delta_{s(C^{\prime},D^{\prime})}$ where $s$ runs over all monomials in $C^{\prime}_{ij},D^{\prime}_{ij}$. Let $V\subset\operatorname{Dist}(G)$ be the $\operatorname{Dist}(G_{*})$-module generated by such $\delta_{s(C^{\prime},D^{\prime})}$. We induct on the degree of $p$ to show that $\delta_{p}$ lies in $V$. The base case, $\deg p=1$, is clear. If $\deg p\geq 1$, write $p=p_{1}(C-1,D)p_{2}(C^{\prime},D^{\prime})$. Recall (e.g. [jantzen_algebraic_groups_2003] 7.7) that for $f_{1},\dots,f_{n}\in I$,
@@ 4.2.5. The Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$
A Steinberg tensor product theorem for $\operatorname{sVec}$ was proved in [kujawa_steinberg_2006] and one for $\operatorname{Ver}_{p}$ was proved in [kannan_steinberg_2024]; we expect to prove Conjecture 4.31 in $\operatorname{Ver}_{4}^{+}$ with similar methods.
[…]

### v4

@@ 3.3. General linear groups and subgroups
where $\underline{\operatorname{End}}$ is an internal Hom. Hence $\mathfrak{gl}(m+nP)$ is an $(m+2n)^{2}$-dimensional vector space, and $d$ acts via $1\otimes d+d\otimes 1$. The bracket operation is given by $\phi_{1},\phi_{2}]=\phi_{1}\phi_{2}-\phi_{2}\phi_{1}+\phi_{2}^{\prime}\phi_{1}^{\prime}$.
@@ 4.2.1. Representations of $\operatorname{GL}(P)$
$0\to\wedge^{3}T_{1}\to T_{1}\otimes\wedge^{2}T_{1}\to S^{(2,1)}T_{1}\to 0.$
@@
For degree $4k$, note that we have a non-split exact sequence $0\rightarrow\chi^{k}\rightarrow T_{4k}\rightarrow\xi\chi^{k}\rightarrow 0$. For degree $4k+1$, note that $T_{4}$ is a nontrivial extension of $\chi$ by $\xi$, so $\chi^{k-1}T_{4}T_{1}$ is a nontrivial extension of $\chi^{k}T_{1}$ by $\xi\chi^{k}T_{1}$. For degree $4k+3$, we showed above that $T_{1}T_{2}$ is a nontrivial extension of $T_{3}$ by $\xi T_{3}$, so $\chi^{k}T_{1}T_{2}$ is a nontrivial extension of $\chi^{k}T_{3}$ by $\xi\chi^{k}\otimes T_{3}$. ∎
@@
Recall that $\mathcal{O}(M_{1})=k[Y,Y^{\prime}]$ with $\Delta(Y)=Y\otimes 1+(1+Y^{\prime})\otimes Y$. Then we can compute $\Delta(Y^{n})$ and $\Delta(Y^{n}Y^{\prime})$ by using that $Y^{4\cdot 2^{k}}$ is primitive and computing $\Delta(Y^{n})$ for $0\leq n<3$, e.g. $\Delta(Y^{2})=1\otimes Y^{2}+Y^{2}\otimes 1+Y^{\prime}\otimes Y^{\prime}$. Note the $Y^{\prime}\otimes Y^{\prime}$ term appears because the braiding is nontrivial, so
@@ 4.2.2. Explicit structure of representations of $\operatorname{GL}(\mathds{1}+P)$ and their tensor products
Let $L(n,T)$ be the irreducible representation of $\operatorname{GL}(1+P)$ corresponding to $n$ an integer (i.e. a representation of $\operatorname{GL}_{1}$) and $T$ an irreducible representation of $\operatorname{GL}(P)$. Because $L(2,\mathds{1})$, $L(0,\chi)$ are one-dimensional, it suffices to describe $L(n,T)$ when $n=0,1$ and $0\leq\deg T<4$.
@@
The weight decompositions for $\operatorname{GL}(\mathds{1}+P)$ representations with highest weight $(n,T)$, $0<\deg T\leq 4$, $0\leq n<2$, are listed below. Extensions are denoted by $\times$.
@@ 4.2.3. Representations of $\operatorname{GL}(nP)$
The augmentation ideal $I\subset\mathcal{O}(\operatorname{GL}(nP))$ is generated by $C_{ij}-\delta_{ij}$ where $\delta$ is the Kronecker delta, and $D_{ij},C^{\prime}_{ij},D^{\prime}_{ij}$ for $1\leq i,j\leq n$. For ease of notation, let $C-1:=\{C_{ij}-\delta_{ij}\}$, $D:=\{D_{ij}\}$, $C^{\prime}:=\{C^{\prime}_{ij}\}$, $D^{\prime}:=\{D^{\prime}_{ij}\}$. Then, the residue classes of monomials in C1C-1, $D$, $C^{\prime}$, $D^{\prime}$ of degree at most $m$ form a basis for $\mathcal{O}(G)/I^{m+1}$ and for a fixed monomial $p(C-1,D,C^{\prime},D^{\prime})$, there is a unique $\delta_{p}\in\operatorname{Dist}(G)$ such that
@@
when $q\neq p$ is another monomial in $C-1,D,C^{\prime},D^{\prime}$ of degree at most $m$. The $\delta_{p}$ over all monomials $p$ of any nonnegative degree form a basis of $\operatorname{Dist}(G)$. We claim that as a $\operatorname{Dist}(G_{*})$-module, $\operatorname{Dist}(G)$ is generated by $\delta_{s(C^{\prime},D^{\prime})}$ where $s$ runs over all monomials in $C^{\prime}_{ij},D^{\prime}_{ij}$. Let $V\subset\operatorname{Dist}(G)$ be the $\operatorname{Dist}(G_{*})$-module generated by such $\delta_{s(C^{\prime},D^{\prime})}$. We induct on the degree of $p$ to show that $\delta_{p}$ lies in $V$. The base case, $\deg p=1$, is clear. If $\deg p\geq 1$, write $p=p_{1}(C-1,D)p_{2}(C^{\prime},D^{\prime})$. Recall (e.g. [20] 7.7) that for $f_{1},\dots,f_{n}\in I$,
@@ 4.2.5. The Steinberg tensor product theorem for $\operatorname{GL}(m+nP)$
A Steinberg tensor product theorem for $\operatorname{sVec}$ was proved in [22] and one for $\operatorname{Ver}_{p}$ was proved in [21]; we expect to prove Conjecture 4.31 in $\operatorname{Ver}_{4}^{+}$ with similar methods.